# Complex Wavelet Mutual Information Loss: A Multi-Scale Loss Function for Semantic Segmentation

### v1

@@ 4.2 Quantitative and qualitative results
As shown in Table 1, the proposed CWMI loss outperforms other loss functions in the majority of metrics for all datasets across U-Net and Attention U-Net models. Furthermore, CWMI consistently outperformed the real CWMI variant, highlighting the importance of incorporating phase information into the steerable pyramid decomposition. Qualitative results from SNEMI3D (Figure 3), GlaS (Figure 4), DRIVE (Figure 5), and MASS ROAD (Figrue 6) further demonstrate the effectiveness of complex CWMI in addressing challenging false positive and false negative segmentation issues that remain unresolved by other state-of-the-art loss functions.
@@ 4.3 Ablation studies
L1 $0.773_{\pm 0.002}$ $0.866_{\pm 0.001}$ $1.22_{\pm 0.05}$ $0.632_{\pm 0.007}$ $0.70_{\pm 0.08}$
L2 $0.774_{\pm 0.003}$ $0.867_{\pm 0.002}$ $1.23_{\pm 0.06}$ $0.633_{\pm 0.011}$ $0.76_{\pm 0.13}$
SSIM $0.468_{\pm 0.032}$ $0.613_{\pm 0.020}$ $5.96_{\pm 0.96}$ $0.081_{\pm 0.030}$ $9.31_{\pm 1.64}$
CWMI $\mathbf{0.779_{\pm 0.004}}$ $\mathbf{0.870_{\pm 0.003}}$ $\mathbf{1.16_{\pm 0.09}}$ $\mathbf{0.640_{\pm 0.005}}$ $\mathbf{0.69_{\pm 0.04}}$
@@
K=2 $0.777_{\pm 0.006}$ $0.868_{\pm 0.004}$ $1.18_{\pm 0.10}$ $0.637_{\pm 0.028}$ $0.79_{\pm 0.02}$
K=4 $\mathbf{0.779_{\pm 0.004}}$ $\mathbf{0.870_{\pm 0.003}}$ $\mathbf{1.16_{\pm 0.09}}$ $\mathbf{0.640_{\pm 0.005}}$ $\mathbf{0.69_{\pm 0.04}}$
K=8 $0.776_{\pm 0.003}$ $0.868_{\pm 0.002}$ $1.20_{\pm 0.06}$ $0.635_{\pm 0.009}$ $0.76_{\pm 0.08}$
K=12 $0.772_{\pm 0.005}$ $0.865_{\pm 0.003}$ $1.33_{\pm 0.05}$ $0.621_{\pm 0.004}$ $0.81_{\pm 0.08}$
     K=4
@@
N=2 $0.760_{\pm 0.002}$ $0.857_{\pm 0.001}$ $1.30_{\pm 0.04}$ $0.618_{\pm 0.009}$ $1.20_{\pm 0.12}$
N=4 $\mathbf{0.777_{\pm 0.002}}$ $\mathbf{0.869_{\pm 0.001}}$ $\mathbf{1.21_{\pm 0.10}}$ $\mathbf{0.634_{\pm 0.017}}$ $\mathbf{0.79_{\pm 0.190}}$
N=6 $0.768_{\pm 0.003}$ $0.862_{\pm 0.002}$ $1.29_{\pm 0.05}$ $0.618_{\pm 0.004}$ $0.88_{\pm 0.18}$
     Layer ablation test
@@
1st layer $0.738_{\pm 0.008}$ $0.843_{\pm 0.006}$ $1.61_{\pm 0.07}$ $0.576_{\pm 0.015}$ $1.38_{\pm 0.14}$
2nd layer $0.762_{\pm 0.008}$ $0.858_{\pm 0.006}$ $1.25_{\pm 0.08}$ $0.624_{\pm 0.016}$ $1.12_{\pm 0.19}$
3rd layer $\mathbf{0.774_{\pm 0.003}}$ $\mathbf{0.867_{\pm 0.002}}$ $\mathbf{1.23_{\pm 0.10}}$ $\mathbf{0.638_{\pm 0.020}}$ $0.82_{\pm 0.12}$
4th layer $0.766_{\pm 0.004}$ $0.861_{\pm 0.003}$ $1.39_{\pm 0.09}$ $0.614_{\pm 0.012}$ $\mathbf{0.76_{\pm 0.04}}$
Table 3: The impact of the decomposition level N and orientation number K on CWMI performance on U-Net with SNEMI3D. In the layer ablation test, only the selected layer is computed for mutual information and all other layers are discard. The bold numbers indicate the best performance for each metric.
@@ 5 Conclusion
In this study, we introduced Complex Wavelet Mutual Information (CWMI) loss, a novel loss function for semantic segmentation that leverages the multiscale, multi-orientation decomposition capabilities of the complex steerable pyramid. By integrating mutual information across wavelet subbands, CWMI effectively captures high-dimensional dependencies and local structural features, including critical phase information, which are essential for accurate segmentation. Extensive experiments on four challenging datasets (SNEMI3D, GlaS, DRIVE, and MASS ROAD) demonstrate that CWMI consistently outperforms state-of-the-art loss functions across most evaluation metrics, particularly in segmenting small instances and narrow boundaries, while introducing minimal computational overhead. These results highlight CWMI as a robust and versatile loss function that effectively addresses key challenges in segmentation, such as class and instance imbalance, boundary precision, and topological consistency.

### v2

@@ 4.2 Quantitative and qualitative results
As shown in Table 1, the proposed CWMI loss outperforms other loss functions across the majority of evaluation metrics for all datasets using both U-Net and Attention U-Net architectures. Statistically, CWMI achieves significantly superior performance on all metrics for the SNEMI3D and DRIVE datasets, as well as on clustering-based metrics (VI and ARI) for the GlaS and MASS ROAD datasets. Moreover, CWMI consistently outperforms its real-valued variant, underscoring the importance of incorporating phase information within the steerable pyramid decomposition. For the Mamba-based architecture, VMUNet, CWMI also significantly outperforms baseline loss functions on the SNEMI3D dataset (Table 2). Qualitative results from SNEMI3D (Figure 3), GlaS (Figure 4), DRIVE (Figure 5), and MASS ROAD (Figure 6) further illustrate CWMI’s effectiveness in mitigating challenging false positive and false negative segmentation errors that persist with other state-of-the-art loss functions.
@@ 4.3 Ablation studies
L1 .$773_{\pm.002}.$$866_{\pm.001}.$ $1.22_{\pm.05}$$632_{\pm.007}.$$70_{\pm.08}.$
L2 .$774_{\pm.003}.$$867_{\pm.002}.$ $1.23_{\pm.06}$$633_{\pm.011}.$$76_{\pm.13}.$
SSIM .$468_{\pm.032}.$$613_{\pm.020}.$ $5.96_{\pm.96}$$081_{\pm.030}.$ $9.31_{\pm 1.64}$
CWMI .$\mathbf{.779_{\pm.004}}$$\mathbf{.870_{\pm.003}}$ $\mathbf{1.16_{\pm.09}}$$\mathbf{.640_{\pm.005}}$$\mathbf{.69_{\pm.04}}$
@@
Contribution of different components of complex steerable pyramid
To better understand the role of individual components in the CWMI loss, we conducted an ablation experiment evaluating the performance of variants based on the mutual information loss of magnitude-only, phase-only, and real-only representations, as shown in Table 4. The CWMI-Phase variant consistently underperformed across all metrics, suggesting that phase information alone is insufficient for reliable segmentation. The CWMI-Mag variant, by contrast, achieved the best performance on clustering metrics (VI and ARI), indicating its strong capacity to capture regional structures. The full CWMI, combining both magnitude and phase via the complex representation, yielded the best results on pixel-wise accuracy metrics (mIoU and mDice) and topological metric (HD), while ranking second on VI and ARI, demonstrating a well-rounded performance across different dimensions of segmentation quality. The CWMI-Real variant performed worse than both CWMI and CWMI-Mag on most metrics, although it outperformed CWMI-Phase, further supporting the necessity of incorporating both magnitude and phase components. These results highlight the critical role of magnitude in structural segmentation, while also demonstrating that the full complex formulation of CWMI achieves superior generalization across diverse evaluation criteria.
[…]
CWMI-Real .$\underline{.776_{\pm.004}}$$\underline{.868_{\pm.003}}$ $1.211_{\pm.048}$$635_{\pm.011}.$$774_{\pm.065}.$
CWMI-Mag .$775_{\pm.004}.$$867_{\pm.003}.$$\mathbf{.908_{\pm.051}}$$\mathbf{.660_{\pm.009}}$$\underline{.735_{\pm.065}}$
CWMI-Phase .$694_{\pm.008}.$$810_{\pm.006}.$ $2.487_{\pm.534}$$458_{\pm.061}.$ $2.000_{\pm.258}$
CWMI .$\mathbf{.778_{\pm.004}}$$\mathbf{.869_{\pm.003}}$ $\underline{1.164_{\pm.079}}$$\underline{.638_{\pm.016}}$$\mathbf{.720_{\pm.066}}$
Table 4: Quantitative comparison of mutual information loss across different components of the complex steerable pyramid decomposition. CWMI-Real: real component; CWMI-Mag: magnitude component; CWMI-Phase: phase component; CWMI: full complex representation (combining real and imaginary components). Results are based on a U-Net model trained on the SNEMI3D dataset. Bold values indicate the best performance, and underlined values denote the second-best for each evaluation metric.
Impact of Regularization Parameter $\lambda$
To evaluate the sensitivity of CWMI to the regularization weight $\lambda$, we experimented with $\lambda=0.1,0.5,0.9$. As shown in Table 5, performance across all evaluation metrics remained largely stable, with only marginal variations. This robustness suggests that CWMI provides complementary information to the cross-entropy term and is not overly sensitive to the regularization strength. Additionally, we found the inclusion of the cross-entropy term to be crucial for directional learning, as mutual information is inherently symmetric and cannot distinguish between correct and inverted label assignments.
@@
$\lambda$=0.1 .$\mathbf{.778_{\pm.004}}$$\mathbf{.869_{\pm.003}}$ $\mathbf{1.188_{\pm.075}}$$636_{\pm.008}.$$\mathbf{.703_{\pm.034}}$
$\lambda$=0.5 .$778_{\pm.006}.$$869_{\pm.004}.$ $1.176_{\pm.076}$$\mathbf{.639_{\pm.014}}$$729_{\pm.078}.$
$\lambda$=0.9 .$777_{\pm.003}.$$869_{\pm.002}.$ $1.219_{\pm.073}$$636_{\pm.013}.$$740_{\pm.061}.$
     $\mathbf{K}$ ablation
@@
K=2 .$777_{\pm.006}.$$868_{\pm.004}.$ $1.18_{\pm.10}$$637_{\pm.028}.$$79_{\pm.02}.$
K=4 .$\mathbf{.779_{\pm.004}}$$\mathbf{.870_{\pm.003}}$ $\mathbf{1.16_{\pm.09}}$$\mathbf{.640_{\pm.005}}$$\mathbf{.69_{\pm.04}}$
K=8 .$776_{\pm.003}.$$868_{\pm.002}.$ $1.20_{\pm.06}$$635_{\pm.009}.$$76_{\pm.08}.$
K=12 .$772_{\pm.005}.$$865_{\pm.003}.$ $1.33_{\pm.05}$$621_{\pm.004}.$$81_{\pm.08}.$
     $\mathbf{N}$ ablation
@@
N=2 .$760_{\pm.002}.$$857_{\pm.001}.$ $1.30_{\pm.04}$$618_{\pm.009}.$ $1.20_{\pm.12}$
N=4 .$\mathbf{.777_{\pm.002}}$$\mathbf{.869_{\pm.001}}$ $\mathbf{1.21_{\pm.10}}$$\mathbf{.634_{\pm.017}}$$\mathbf{.79_{\pm.190}}$
N=6 .$768_{\pm.003}.$$862_{\pm.002}.$ $1.29_{\pm.05}$$618_{\pm.004}.$$88_{\pm.18}.$
     Layer ablation
[…]
1st layer .$738_{\pm.008}.$$843_{\pm.006}.$ $1.61_{\pm.07}$$576_{\pm.015}.$ $1.38_{\pm.14}$
2nd layer .$762_{\pm.008}.$$858_{\pm.006}.$ $1.25_{\pm.08}$$624_{\pm.016}.$ $1.12_{\pm.19}$
3rd layer .$\mathbf{.774_{\pm.003}}$$\mathbf{.867_{\pm.002}}$ $\mathbf{1.23_{\pm.10}}$$\mathbf{.638_{\pm.020}}$$82_{\pm.12}.$
4th layer .$766_{\pm.004}.$$861_{\pm.003}.$ $1.39_{\pm.09}$$614_{\pm.012}.$$\mathbf{.76_{\pm.04}}$
Table 5: The impact of regularization parameter $\lambda$, decomposition level N, and orientation number K on CWMI performance on U-Net with SNEMI3D. In the layer ablation test, only the selected layer is computed for mutual information and all other layers are discard. The bold numbers indicate the best performance for each metric.
@@ 5 Conclusion
In this study, we introduced Complex Wavelet Mutual Information (CWMI) loss, a novel loss function for semantic segmentation that leverages the multiscale, multi-orientation decomposition capabilities of the complex steerable pyramid. By integrating mutual information across wavelet subbands, CWMI effectively captures high-dimensional dependencies and local structural features, including critical phase information, which are essential for accurate segmentation. Extensive experiments on four challenging datasets demonstrate that CWMI consistently outperforms state-of-the-art loss functions across most evaluation metrics, particularly in segmenting small instances and narrow boundaries, while introducing minimal computational overhead. These results highlight CWMI as a robust and versatile loss function that effectively addresses key challenges in segmentation, such as class and instance imbalance, boundary precision, and topological consistency.
@@
Acknowledgements
This study was conducted without external funding. I would like to thank the reviewers for their constructive feedback, which significantly improved the quality of this work. I am especially grateful to my wife, Di Duan, for her unwavering support of my passion for machine learning. I thank Xinzi He for valuable discussions that contributed to the revision of this study.